NASA is expected to send a 2600-kg satellite 450 km above the earth’s surface.


(a) What is its radius?


(b) What speed will it have?


(c) What is its orbital period?


(d) What is its radial acceleration?

1
Expert's answer
2021-11-14T17:12:03-0500

(a) We can find the radius of the satellite as follows:


"R=R_E+h=6.37\times10^6\ m+4.5\times10^5\ m=6.82\times10^6\ m."

(b) We can find the speed of the satellite as follows:


"F_c=F_g,""\dfrac{mv^2}{R}=\dfrac{GM_Em}{R^2},""v=\sqrt{\dfrac{GM_E}{R}}=\sqrt{\dfrac{6.67\times10^{-11}\ \dfrac{N\times m^2}{kg^2}\times5.98\times10^{24}\ kg}{6.82\times10^6\ m}},""v=7647\ \dfrac{m}{s}."

(c) We can find the orbital period of the satellite as follows:


"v=\dfrac{2\pi R}{T},""T=\dfrac{2\pi R}{v}=\dfrac{2\pi\times6.82\times10^6\ m}{7647\ \dfrac{m}{s}}=5604\ s,""T=5604\ s\times\dfrac{1\ h}{3600\ s}=1.56\ h."

(d) We can find the radial (centripetal) acceleration as follows:

"a_c=\dfrac{v^2}{R}=\dfrac{(7647\ \dfrac{m}{s})^2}{6.82\times10^6\ m}=8.57\ \dfrac{m}{s^2}."